OpenStudy (anonymous):

The velocity of sound in air is given by the equation v=20r273+t where v is the velocity in meters per second and t is the temperature in degrees Celsius. Find the velocity when the temperature is 117ºC. Round to the nearest meter/sec. A. 20 meters per second B. 330 meters per second C. 395 meters per second D. 216 meters per second

OpenStudy (ganpat):

what r in between of equation ??

OpenStudy (anonymous):

square root

OpenStudy (ganpat):

assuming it ok.. So Velocity = \[20 * \sqrt{273} + t = 20 * \sqrt{273} + 117 = \] = 330.45 +117 it comes out to be 447 m/s...

OpenStudy (turingtest):

@chadandtricia_2009 better to write v=20sqrt(273)+t it makes it more clear for us to read, and the parentheses make it clear where the root ends thanks

OpenStudy (turingtest):

because I'm not sure whether you mean v=20sqrt(273)+t or v=20sqrt(273+t) see the problem?

OpenStudy (anonymous):

ok sorry. can you help me with this one please?

OpenStudy (turingtest):

well which is it? first or second one?

OpenStudy (anonymous):

second

OpenStudy (turingtest):

put the temperature given, 117, where the t is in your equation then use a calculator

OpenStudy (anonymous):

273+117=390

OpenStudy (anonymous):

20sqrt390

OpenStudy (turingtest):

yup

OpenStudy (anonymous):

so then what do I do?

OpenStudy (turingtest):

again, please use parentheses 20sqrt(390) it's just easier on the eyes now? simplify if possible, which I don't think it is to see if it can be simplified find the prime factorization of the number under the radical, 390

OpenStudy (anonymous):

no it can't. I think the answer is c. 395

OpenStudy (turingtest):

yes :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

395 is my answer